OpenStudy (calculusxy):

MEDAL! Maryam used an exercise bicycle on Tuesday and Friday. On Tuesday, she bicycled 15.25 miles per hour for 40 minutes and an additional 7.5 miles per hour, for 30 minutes. On Friday she bicycled a total of 12.5 miles. How many more miles did Maryam bicycle on Tuesday than on Friday?

OpenStudy (calculusxy):

@amistre64

OpenStudy (calculusxy):

@ganeshie8

OpenStudy (amistre64):

add up the miles

OpenStudy (calculusxy):

22.75

OpenStudy (amistre64):

15.25 (40/60) + 7.5 (30/60)

OpenStudy (anonymous):

The first day Tuesday

OpenStudy (amistre64):

rate per hour times portion of hours

OpenStudy (amistre64):

spose you are going 1 mph and ride for 98 minutes youve gone 98/60 miles

OpenStudy (anonymous):

Because 15.25 mph and 40 minutes that's 2/3 of 15.25 so that would mean 10.1666666 and so on. That plus 30 minutes of 7.5 mph which is 1/2 of 7.5 is 3.75 that = 13. Something and that is more than Friday

OpenStudy (calculusxy):

So I did that and got 167/12 which would equal to 13 11/12 miles. Now do I subtract that with 12.5 (12 1/2)

OpenStudy (amistre64):

by subtracting of course

OpenStudy (amistre64):

1/2 = 6/12 right?

OpenStudy (calculusxy):

Yeah

OpenStudy (amistre64):

then 13 - 12 and (11-6)/12

OpenStudy (anonymous):

hinduism ki jai

OpenStudy (anonymous):

me hindu too

OpenStudy (calculusxy):

My answer was 13 5/12 right?

OpenStudy (anonymous):

amistre please feel mercy at me

OpenStudy (calculusxy):

@Math.Genius Hare Krishna

OpenStudy (amistre64):

not quite 13 + 11/12 -12 - 6/12 ------------ 1 + 5/12

OpenStudy (anonymous):

hare rama hare krishna krishna krishna hare hare

OpenStudy (anonymous):

bharat mata ki jai

OpenStudy (amistre64):

Math, you have, or had, someone capable assisting you. if you want to do well with the test, then learn from them. i was unable to interpret your questions approriately.

OpenStudy (calculusxy):

@amistre64